Three electric bulbs of rating 60 W are joined in series and then connected to electric mains . The power consumed will be ?

Grade:10

1 Answers

Karthik Kumar
8 years ago
Let resistance of each bulb is R
then total resistance of 3 bulbs=3R
power=V2/R  =60W for each bulb
when three bulbs connected in series  then power=
V2/(3R)
=(1/3)(V2/R)
=(1/3)(60)
=60/3 =20 W
(voltage is same when one bulb is used or three used)